OpenStudy (anonymous):

A woman went marketing for sugar for canning. In one store sugar was 8 1/2 cents per pound and she found that she lacked 30 cents of having enough money to buy the number of pounds she needed. She bought the required amount of sugar at another store at 8 cents per pound, and had 45 cents left. How much money did she have before making the purchase?

OpenStudy (anonymous):

where did this problem come from? a book from 1943?

OpenStudy (anonymous):

I'm pretty sure it was o_o

OpenStudy (anonymous):

i bet we can do it, if we think hard enough need to some up with two equations and two unknowns i think one is the amount of money trixie has, lets call that \(x\) and the other, the amount of sugar she needs, lets call it \(y\)

OpenStudy (anonymous):

the first line In one store sugar was 8 1/2 cents per pound and she found that she lacked 30 cents of having enough money to buy the number of pounds she needed. can be interpreted as \(8.5y=x+30\)

OpenStudy (anonymous):

Are there answer choices? U can plug them into an equation, and solve.

OpenStudy (anonymous):

the second line She bought the required amount of sugar at another store at 8 cents per pound, and had 45 cents left. can be interpreted as \[8y=x-45\]

OpenStudy (anonymous):

got any idea about how to solve those?

OpenStudy (anonymous):

No, sorry.

OpenStudy (anonymous):

since it is the same \(x\) in both we can write them as \[x=8.5y-30\\ x=8y+45\]

OpenStudy (anonymous):

that means \[8.5y-30=8y+45\] can you solve that one?

OpenStudy (anonymous):

y= 150

OpenStudy (anonymous):

yes then solve for \(x\)

OpenStudy (anonymous):

you can plug it in here \[x=8y+45\\ x=8\times 75+45\]

OpenStudy (anonymous):

x= 1245